satify the equation of the lines
  • cos2x-sinx+5=0
  • 4cosx+secx-4=0

Grade:11

1 Answers

Xyz
3 years ago
1- 2 sin^2 X - sinx +5=0.   (We know that cos2x= 1-2sin^2x)
2sin^2 x+sinx -6=0
(2sinx-3)(sinx+2)=0
We got two value of sinx= 3/2, -2
From second equation 4cos^2 X - 4cosx +1 =0. (we know that secx = 1/cosx
(2cox-1)^2=0
We got cosx= ½
It means these two equations does not cross each other at any point
Thanks and keep asking if any doubt